# Automatic Outlier Rectification via Optimal Transport

### v1

@@ C.1 Mean Estimation Details
C.2 Concave Cost Mean Estimation Simulation
@@
C.3 Convex Cost Mean Estimation Simulation
@@
C.4 Least Absolute Regression Details
The experimental details for least absolute regression follow. For each $(\delta,r)$ point in our experiments, we perform 100 random trials at 100 fixed seeds. In each trial, we run our optimization procedure for with a learning rate of $10^{-2}$ starting from 10 randomly initialized points and take the best loss. In each problem, we stop optimization when the number of iterations reaches 1000 or the change in the loss function between successive iterations is below a tolerance of $10^{-6}$. We initialize the slope $\beta$ according to a $N(0,1)$ distribution and the bias to zero.
We provide additional results on the least absolute deviation simulations below. In Table 4 we report the results of our experiment. We also test our cross-validation procedure in the LAD regression context to show that it improves the selection of the budget $\delta$. We take the same experimental setting as in the regression example from the main text. In particular, our algorithm for cross-validation which proceeds as follows:
  1. 1.
    First, we draw a small sample of size $N=8$ from the dataset. If a corruption level is suspected, it may be used to find a reasonable sample size for this purpose using a binomial probability calculation. We have tried several small values from $N=5$ to $N=10$ and found the performance to be relatively similar.
  2. 2.
    Using this sample, we fit our robust estimator for a range of values of $\delta$, the budget parameter. We use $\delta\in\{0.5,1.0,1.5,2.0,2.5,3.0\}$. We select the $\delta$ which produces the smallest loss on this sample data set.
  3. 3.
    Using the optimal $\delta$ selected in step 2, we fit our robust estimator on the entire dataset. The results of this procedure are displayed in Table 4.
The results are included in Table 4. As can be seen, our estimator outperforms all competing estimators, and attains significant outperformance for most corruption levels. Cross-validation significantly improves the results for several settings of $\epsilon$, but performs worse than our fixed $\delta$ for some of the larger values of $\epsilon$. We hypothesize that this is due to the impossibility of distinguishing which distribution is the “correct” distribution as $\epsilon\to 0.5$, which may lead to learning regression slopes which fit the corruption distribution instead of the clean distribution. However, for smaller values of $\epsilon$, we see a moderate improvement due to cross-validation.
@@ C.4 Least Absolute Regression Details
Ours 0.803 $\pm$ 0.860 0.574 $\pm$ 0.778 0.355 $\pm$ 0.878 0.462 $\pm$ 0.962 0.572 $\pm$ 0.998
Ours (CV) 0.349 $\pm$ 0.784 0.509 $\pm$ 0.898 0.667 $\pm$ 1.184 0.832 $\pm$ 1.354 0.750 $\pm$ 1.282
Table 4: We compared our estimator with several standard regression methods by evaluating the average loss on clean data points across various corruption levels. Mean performance is accompanied by a 95% confidence interval over 100 random trials. In our evaluation, we set the threshold parameter of Huber regression to 1.5. The hyperparameters for our estimator, namely $\delta=1.5$ and $r=0.5$, remained constant across all corrupted levels.
@@
C.5 Concave Cost Regression Simulation
@@
C.6 Convex Cost Regression Simulation
@@
C.7 Additional Volatility Surface Experiment
@@
Our cross-validation procedure splits the train day into a training and validation sample, which we use to obtain estimates of MAPE and $\nabla\hat{S}$ as a function of $\delta$. We select the delta that optimizes MAPE and $\nabla\hat{S}$ and fit on the entire train day. Then, on the out-of-sample test day, we evaluate our method for the optimally selected $\delta$. Thus, to be clear, for the estimated optimal $\delta$, we estimate the surface on one day, and evaluate it completely out-of-sample on the consecutive trading day.
The results are collected in the table below. The results of our experiment show that our cross-validation approach outperforms competing approaches and improves upon the fixed $\delta$ results reported in the main text. The below tables show the out-of-sample MAPE and $\nabla\hat{S}$ averaged over all out-of-sample test days.
@@
C.8 Options Volatility Surface Details
@@
We implement the KS estimator and our estimator in PyTorch. For our estimator, we estimate the surface by subgradient descent with learning rate $\alpha=10^{-1}$ and $r=0.5$, terminating when the relative change in loss reaches $10^{-5}$. We denote the test data set’s option’s implied volatilities $y^{\prime}_{i}$ and the estimated surface’s implied volatility for option $i$ as ${\hat{S}}(x^{\prime}_{i})$. We begin each iteration by setting $\delta=\ell(x,y)/2\|theta\|^{r}$, where $\ell$ is the loss function of Theorem 2. We initialize $\theta$ to those of the benchmark estimator. We perform 5 trials per options chain. In each trial, we randomly select a different 80% train and 20% test set, estimate the surface on the train set, and record the two losses (MAPE and $\nabla\hat{S}$) on the test set. We depict all losses gathered in this way via the histograms of Section 6. Experiments are run on a server with a Xeon E5-2398 v3 processor and 756GB of RAM.
@@
The cross-validation (CV) procedure samples 2/3 of the training surface as a CV training set and leaves the remaining 1/3 as the CV validation set. Empirically, we expect $\epsilon$ to be quite small (e.g. $\epsilon\in(0,0.05]$) for this data set, which allows us to take a larger cross-validation set, as most option chains contain only a few hundred options. Five such random splits are drawn and used to estimate the MAPE of $\delta\in\{10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1},1,2,5,10\}$. The MAPEs of each $\delta$ are averaged across the five CV runs and the $\delta$ with the lowest MAPE is then used to fit our estimator on the entire option chain.
@@
We define the option implied volatility surface $\hat{S}(x)$ as the function from option feature vectors $x$ to estimated implied volatilities $y$. That is, for some given $x$, we have $y:=\hat{S}(x)$. The accuracy measure MAPE is defined as follows for some test set of options $\{(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{n}$:
@@
where $c$ is a small numerical stability factor we set to $0.01$. This is a suitable choice, as $\geq$99of implied volatilities are greater than 0.1.
@@
In the expression above, $\hat{S}_{\tau_{i},\Delta_{j}}:=\hat{S}(\log\tau_{i},u(\Delta_{j}),z)$ where $z=1$ if $\Delta_{j}>0$ and 0 otherwise. Here, each tuple ($\tau_{i}$, $\Delta_{j}$) is a member of the Cartesian product of $\{\tau_{j}:j\in[p]\}$ and $\{\Delta_{k}:k\in[s]\}$, where ”$[m]$” denotes the set of natural numbers from 0 to $m$, and $p=10$ and $s=39$. By convention, OptionMetrics selects the following 11 discretization points for $\tau$, which we follow:
@@
Figure 15: This plot depicts the option implied volatility surface $\hat{S}(x)$ estimated by our estimator with $\delta=1$ on the displayed options, which are depicted as blue and red dots. Blue denotes call options and red denotes put options. Outliers are present in this option chain near 0-100 days until expiration for deltas around -0.2 and 1.0. However, these outliers do not corrupt the fitted surface, which fits options nearby the outlying options quite well as compared to the surface fit by the benchmark method displayed in the figure above. Additionally, in contrast to the surface fit by the benchmark KS method, the surface fit by our estimator only reaches values of $\sim 65$% annualized implied volatility and has a surface gradient $\nabla\hat{S}$ of 0.69, a smoothness improvement of over 100%. This example illustrates the capability of our estimator to rectify outliers.

### v2

@@ F.2 Mean Estimation Details
F.3 Concave Cost Mean Estimation Simulation
@@
F.4 Convex Cost Mean Estimation Simulation
@@
F.5 Least Absolute Regression Details
We provide additional results on the least absolute deviation simulations below. In Table 4 we report the results of our experiment. We take the same experimental setting as in the regression example from the main text.
For each $(\delta,r)$ point in our experiments, we perform 100 random trials at 100 fixed seeds. In each trial, we run our optimization procedure for with a learning rate of $10^{-2}$ starting from 10 randomly initialized points and take the best loss. In each problem, we stop optimization when the number of iterations reaches 1000 or the change in the loss function between successive iterations is below a tolerance of $10^{-6}$. We initialize the slope $\beta$ according to a $N(0,1)$ distribution and the bias to zero.
The results are included in Table 4. As can be seen, our estimator outperforms all competing estimators, and attains significant outperformance for most corruption levels.
Table 4: We compared our estimator with several standard regression methods by evaluating the average loss on clean data points across various corruption levels. Mean performance is accompanied by a 95% confidence interval over 100 random trials. In our evaluation, we set the threshold parameter of Huber regression to 1.5. The hyperparameters for our estimator, namely $\delta=1.5$ and $r=0.5$, remained constant across all corrupted levels. Error bars represent two standard deviation confidence intervals we have computed manually assuming normal errors.
@@ F.5 Least Absolute Regression Details
Ours 0.657 $\pm$ 0.638 0.529 $\pm$ 0.430 0.680 $\pm$ 0.467 0.802 $\pm$ 0.597 0.866 $\pm$ 0.619
@@
F.6 Concave Cost Regression Simulation
@@
F.7 Convex Cost Regression Simulation
@@
Appendix G Additional Volatility Modeling Details & Experiments
G.1 Additional Volatility Surface Experiment
@@
To perform cross-validation, we split the train day into a training and validation sample, which we use to obtain estimates of MAPE and $\nabla\hat{S}$ as a function of $\delta$. We select the delta that optimizes MAPE and $\nabla\hat{S}$ and fit on the entire train day. Then, on the out-of-sample test day, we evaluate our method for the optimally selected $\delta$. Thus, to be clear, for the estimated optimal $\delta$, we estimate the surface on one day, and evaluate it completely out-of-sample on the consecutive trading day.
The results are collected in the table below. The results of our experiment show that our approach outperforms competing approaches and that cross-validation improves upon the fixed $\delta$ results reported in the main text. The below tables show the out-of-sample MAPE and $\nabla\hat{S}$ averaged over all out-of-sample test days.
@@
G.2 Kernelized Regression Experiment Details
@@
We implement the KS estimator and our estimator in PyTorch. For the 2SKS method, options in an options chain are removed if their implied volatilities fall outside of the region $[q_{0.25}-1.5\cdot IQR,q_{0.75}+1.5\cdot IQR]$, where $q_{0.25}$ is the 25% quantile of the implied volatilities in the options chain, $q_{0.75}$ is defined similarly, and IQR is the interquartile range of the implied volatilities in the options chain. The surface is then estimated upon the remaining options via the KS method. For our estimator, we estimate the surface by subgradient descent with learning rate $\alpha=10^{-1}$ and $r=0.5$, terminating when the relative change in loss reaches $10^{-5}$. We denote the test data set’s option’s implied volatilities $y^{\prime}_{i}$ and the estimated surface’s implied volatility for option $i$ as $\hat{S}(x^{\prime}_{i})$. We begin each iteration by setting $\delta=\ell(x,y)/2\norm{\theta}^{r}$, where $\ell$ is the loss function of Theorem 4. We initialize $\theta$ to those of the benchmark estimator. We perform 5 trials per options chain. In each trial, we randomly select a different 80% train and 20% test set, estimate the surface on the train set, and record the two losses (MAPE and $\nabla\hat{S}$) on the test set. We depict all losses gathered in this way via the histograms of Section 5.2. Experiments are run on a server with a Xeon E5-2398 v3 processor and 756GB of RAM.
@@
The cross-validation (CV) procedure is standard. We sample 4/5 of the training surface as a CV training set and leave the remaining 1/5 as the CV validation set. Five such splits are made and used to estimate the MAPE of $\delta\in\{10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1},1,2,5,10\}$. The MAPEs of each $\delta$ are averaged across the five CV runs and the $\delta$ with the lowest MAPE is then used to fit our estimator on the entire option chain.
@@
These losses are chosen for their importance in finance and option trading and valuation. MAPE is a preferable error metric for volatility surfaces, as option chains contain options with implied volatilities which can differ in orders of magnitude, and MAPE weighs equally the contribution to error of options with such volatilities. The discretized surface gradient $\nabla\hat{S}$ measures surface smoothness of the estimated volatility surface by computing a discrete gradient across a grid of points in the (time to exporation, strike price) plane. Smooth surfaces are important for several reasons: (1) smoother surfaces allow more stable interpolation or extrapolation of implied volatility for options with market prices which are not directly observable; (2) smoother surfaces produce smaller adjustments for option hedging positions constructed from measures of the IVS, which ultimately lowers hedging transaction costs; (3) smoother surfaces are less likely to have internal arbitrages between options, which are implied by sharp discontinuities in the IVS, putting the surface more in line with established asset pricing theory; and (4) smoother surfaces produce more stable estimates of financial institutions’ derivative exposures, which are required to be consistent for financial regulatory and reporting requirements. These measures are standard in options IVS estimation. We define the option implied volatility surface $\hat{S}(x)$ as the function from option feature vectors $x$ to estimated implied volatilities $y$. That is, for some given $x$, we have $y:=\hat{S}(x)$. The accuracy measure MAPE is defined as follows for some test set of options $\{(x^{\prime}_{i},y^{\prime}_{i})\}_{i=1}^{n}$:
@@
where $c$ is a small numerical stability factor we set to $0.01$. This is a suitable choice, as $\geqslant$99of implied volatilities are greater than 0.1.
@@
In the expression above, $\hat{S}_{\tau_{i},\Delta_{j}}:=\hat{S}(\log\tau_{i},u(\Delta_{j}),z)$ where $z=1$ if $\Delta_{j}>0$ and 0 otherwise. Here, each tuple ($\tau_{i}$, $\Delta_{j}$) is a member of the Cartesian product of $\{\tau_{j}:j\in[p]\}$ and $\{\Delta_{k}:k\in[s]\}$, where "$[m]$" denotes the set of natural numbers from 0 to $m$, and $p=10$ and $s=39$. By convention, OptionMetrics selects the following 11 discretization points for $\tau$, which we follow:
@@
Figure 16: This plot depicts the option implied volatility surface $\hat{S}(x)$ estimated by our estimator with $\delta=1$ on the displayed options, which are depicted as blue and red dots. Blue denotes call options and red denotes put options. Outliers are present in this option chain near 0-100 days until expiration for deltas around -0.2 and 1.0. However, these outliers do not corrupt the fitted surface, which fits options nearby the outlying options quite well as compared to the surface fit by the benchmark method displayed in Figure 15. Additionally, in contrast to the surface fit by the benchmark KS method, the surface fit by our estimator only reaches values of $\sim 65$% annualized implied volatility and has a surface gradient $\nabla\hat{S}$ of 0.69, a smoothness improvement of over 100%. This example illustrates the capability of our estimator to rectify outliers.
G.3 Deep Learning Experiment Details
The full details and description of the deep learning local volatility surface estimation experiment are given in this section.
Background.
IVS estimation provides an approximate surface fit for the volatilities implied by option market prices. However, market participants often require surfaces which fit as closely as possible to the market-implied volatilities. [^1] For example, banks pricing some kinds of exotic options require such surfaces to avoid arbitrage opportunities in surfaces that their internal trading teams may exploit, as outlined in Hull and Basu [2022]. Participants pricing options may moreover simply prefer to accept the assumption that the market prices are usually correct, and thus prefer closely fitting surfaces. One of the most popular methods used in mathematical finance to estimate surfaces obeying such a constraint under a reasonable stochastic process model for the evolution of asset prices is the local volatility or implied volatility function model introduced by Rubinstein [1994], Dupire et al. [1994], and Emanuel Derman and Zou [1996]. The local volatility model assumptions imply that a function $\sigma(K,T)$ exists which gives an analytic expression for the volatility $\sigma$ at a given option strike price $K$ and option time to expiration $T$. This function is known as Dupire’s formula after its originator. Enforcing Dupire’s formula in the estimation of a surface results in a local volatility surface, which is of great use to financial market participants. However, estimating such a surface becomes significantly complicated if market data contains corruptions or unrealistic outliers which are unlikely to occur in the future. [^2] This may occur due to low liquidity, which allows market manipulation or adversarial trading to produce unrealistic prices. As human intervention for outlier rectification in a large number of IVS fitting tasks is unrealistic, an automatic method for rectifying outliers in models fitting local volatility surfaces is required.
Data.
We utilize the data set of Chataigner et al. [2020], which contains 17 (options chain, IVS) pairs taken from the German DAX index in August 2001. Options chains are captured from daily options market data, and IVSes are estimated for a grid of strikes and maturities via a trinomial tree calibrated by the method of Crépey [2002]. In addition to Chataigner et al. [2020], this data set is also used in Crépey [2002] and Crépey [2004]. In this data set’s usage in the literature, the authors train and test using the price surface, which is given by a nonlinear transformation of the IVS. Consequently, we also use the price surface for fitting; however, we report results after transforming back to the IVS. To corrupt the data set, a percentage of the options in each options chain are selected. We refer to this percentage, the corruption level, as $\varepsilon$. The prices of these options $x$ are then corrupted by replacing them with the values $10x$.
Benchmark Method.
As a benchmark, we apply the state-of-the-art deep learning approach from Chataigner et al. [2020], the Dupire neural network. In this approach, a neural network estimates the price surface under the local volatility model assumptions encoded in Dupire’s formula, and additional no-arbitrage constraints which are useful for empirical applications. These conditions are enforced using different approaches: hard, split, and soft constraints. Hard constraints enforce the conditions by separating the network into subnetworks which have separate input layers for the variables involved in each of the conditions; conditions such as convexity and non-negativity are then enforced by projection steps during optimization and the proper choice of activation functions. Hard constraints thus learn a function which explicitly satisfies the conditions. Soft constraints enforce these conditions instead by penalizing violations of them at each of the training points observed from the market. Hybrid constraints split the network into subnetworks which have separate input layers for each of the constrained variables as in hard constraints, and utilize soft constraints for enforcing the Dupire and no-arbitrage conditions. The function learned only approximately satisfies the conditions in theory, but in practice the approximation is very good (see Chataigner et al. [2020] for details). Because these approaches estimate the price surface, we also estimate the price surface in this application; however, the local volatility surface is easily recovered from the price surface by a nonlinear transformation.
Robust Neural Network Estimation Problem & Implementation.
To estimate price and volatility surfaces under the induced corruption, we estimated the benchmark approach using our statistically robust estimator. For each options chain, we tune $\delta$ via cross-validation: for each day, we sample 80% of the training set without replacement as a cross-validation (CV) training set, and use the remaining 20% of the training set as a CV validation set. We then train our robust neural network estimator on the CV training set with $\delta\in\{10^{-3},10^{-2},...,10^{3}\}$. For each $\delta$, we then compute the MAPE on the CV validation set. We then select the $\delta$ with the lowest MAPE on CV validation set, and we use this $\delta$ to train our estimator on the entire training set. This produces our final estimated model. We base our implementation on the implementation of Chataigner et al. [2020] (see commit e03da98 at this url), which is BSD 3-clause licensed. The experiments were run on a server with a Xeon E5-2398 v3 processor with 756GB of RAM.
Results.
To maintain comparability with the benchmark, we test and report the same metrics reported by Chataigner et al. [2020], namely, the RMSE and MAPE for the surface estimation. The metrics for each day and model are repeated for three trials with different random weight initialization seeds. The results are displayed in Table 2. The first panel displays the RMSE and MAPE for each model using the different constraint techniques, averaged across all trials and corruption levels. The second panel displays the RMSE and MAPE for each corruption level averaged across all models. Our approach outperforms the baseline approach across all averages, and does so more clearly as the corruption level increases, despite the strong regularizing effect of the no-arbitrage constraints and enforcement of Dupire’s formula. Our improvement is present across all models, and is especially impactful for the most accurate model utilizing soft arbitrage constraints. For this model, the out-of-sample test error in terms of RMSE and MAPE are reduced by 33% and 34%, respectively. This experiment displays the efficacy of our estimator in a state-of-the-art deep learning model for option price surface modeling.